(x-7)/3 = (x-1)/5
(x-7)/3 = (x-1)/5(x-7)/3 = (x-1)/5(x-7)/3 = (x-1)/5(x-7)/3 = (x-1)/5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(x-7)/3 = (x-1)/5

5(x-7) = 3(x-1)

5x – 35 = 3x – 3

5x – 3x = -3 + 35

2x = 32

x = 16

Thanks